# Thermodynamic Variational Principle Unifying Gravity and Heat Flow

### v1

@@ Abstract
Let $\bar{p}^{\mathrm{\ell}}$ and $\bar{p}^{\mathrm{u}}$ be the mean pressures in regions $\mathrm{\ell}$ and $\mathrm{u}$, respectively. The first equation in (14) gives
$(\bar{p}^{\mathrm{\ell}}-\bar{p}^{\mathrm{u}})V=\frac{NmgL}{2}.$ (A1)
Assuming linear pressure profiles in each phases $\mathrm{\ell}$ and $\mathrm{u}$, this relation reflects the overall force balance $(p_{1}-p_{2})A=Nmg$, and implies pressure continuity at the interface.
The second equation of (14) yields
$\mu^{\mathrm{\ell}}-\mu^{\mathrm{u}}-(S^{\mathrm{\ell}}+S^{\mathrm{u}})\frac{\Xi}{2N}=\frac{mgL}{2}.$ (A2)
Here $\mu^{\mathrm{\ell}}=\mu^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},\bar{p}^{\mathrm{\ell}})$ and $\mu^{\mathrm{u}}=\mu^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},\bar{p}^{\mathrm{u}})$ are the chemical potentials of phases $\mathrm{\ell}$ and $\mathrm{u}$. Using $T_{\rm s}$ defined in Eq. (16) of the main text, the global temperatures for each phase, given by Eq. (13), can be written as
$\tilde{T}^{\mathrm{\ell}}=T_{\rm s}-\frac{\Xi}{2}\frac{{\mathcal{N}}^{\ell}}{N},\quad\tilde{T}^{\mathrm{u}}=T_{\rm s}+\frac{\Xi}{2}\frac{N-{\mathcal{N}}^{\ell}}{N}.$ (A3)
Furthermore, the mean pressures $\bar{p}^{\mathrm{\ell}}$ and $\bar{p}^{\mathrm{u}}$ relate to the interface pressure $p_{\theta}$ via the force balance relations for each phase:
$(\bar{p}^{\mathrm{\ell}}-p_{\theta})A=\frac{{\mathcal{N}}^{\ell}mg}{2},\quad(\bar{p}^{\mathrm{u}}-p_{\theta})A=-\frac{(N-{\mathcal{N}}^{\ell})mg}{2}.$ (A4)
Expanding $\mu^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},\bar{p}^{\mathrm{\ell}})$ and $\mu^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},\bar{p}^{\mathrm{u}})$ around $(T_{\rm s},p_{\theta})$ up to $O(\varepsilon)$, using $\partial\mu/\partial T=-S/N$, $\partial\mu/\partial p=V/N$, ${\mathcal{V}}^{\ell}=Ax_{\theta}$, and $V-{\mathcal{V}}^{\ell}=A(L-x_{\theta})$ along with (A3) and (A4), we have
$\mu^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},\bar{p}^{\mathrm{\ell}})=\mu^{\mathrm{\ell}}(T_{\rm s},p_{\theta})+\frac{\Xi}{2}\frac{S^{\mathrm{\ell}}}{N}+\frac{mgx_{\theta}}{2}+O(\varepsilon^{2}),$ (A5)
$\mu^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},\bar{p}^{\mathrm{u}})=\mu^{\mathrm{u}}(T_{\rm s},p_{\theta})-\frac{\Xi}{2}\frac{S^{\mathrm{u}}}{N}-\frac{mg(L-x_{\theta})}{2}+O(\varepsilon^{2}).$ (A6)
Substituting (A5) and (A6) into (A2), and neglecting terms of $O(\varepsilon^{2})$, we obtain Eq. (15) of the main text: $\mu^{\mathrm{\ell}}(T_{\rm s},p_{\theta})=\mu^{\mathrm{u}}(T_{\rm s},p_{\theta})$.
@@
Consider the solution $(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})$ where the liquid phase settles in the lower region. We write
$N^{\rm L}_{\rm s}=N^{\rm L}_{0}+\Delta N,\quad V^{\rm L}_{\rm s}=V^{\rm L}_{0}+\Delta V,$ (B7)
where $N^{\rm L}_{0}$ and $V^{\rm L}_{0}$ are the particle number and volume of the liquid phase at $g=0$ and $\Xi=0$ (determined by the lever rule), and $\Delta N,\Delta V$ are deviations of $O(\varepsilon)$. According to Eq. (12) in the main text, the variational function is
${\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})=F^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},V^{\rm L}_{\rm s},N^{\rm L}_{\rm s})$
$~{}+F^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},V-V^{\rm L}_{\rm s},N-N^{\rm L}_{\rm s})-\Psi(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})mgL.$ (B8)
@@
$\Psi(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})=\frac{N}{2}\left(\frac{N^{\rm L}_{0}}{N}-\frac{V^{\rm L}_{0}}{V}\right)+O(\varepsilon),$ (B9)
$\tilde{T}^{\mathrm{\ell}}=\tilde{T}-\frac{\Xi}{2}\frac{N^{\rm G}_{0}}{N}+O(\varepsilon^{2}),$ (B10)
$\tilde{T}^{\mathrm{u}}=\tilde{T}+\frac{\Xi}{2}\frac{N^{\rm L}_{0}}{N}+O(\varepsilon^{2}).$ (B11)
@@
$F^{\mathrm{\ell}}(\tilde{T}^{\mathrm{\ell}},V^{\rm L}_{\rm s},N^{\rm L}_{\rm s})+F^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},V-V^{\rm L}_{\rm s},N-N^{\rm L}_{\rm s})$
$=F(\tilde{T},V^{\rm L}_{0},N^{\rm L}_{0})+F(\tilde{T},V^{\rm G}_{0},N^{\rm G}_{0})$
$\quad-S_{0}^{\mathrm{L}}(\tilde{T}^{\mathrm{\ell}}-\tilde{T})-p_{0}^{\mathrm{L}}\Delta V+\mu_{0}^{\mathrm{L}}\Delta N$
@@
Here, the subscript ‘0’ indicates the quantity at zero-gravity in equilibrium. Using the equilibrium conditions at $O(\varepsilon^{0})$, $p_{0}^{\mathrm{L}}=p_{0}^{\mathrm{G}}$ and $\mu_{0}^{\mathrm{L}}=\mu_{0}^{\mathrm{G}}$, the terms proportional to $\Delta V$ and $\Delta N$ cancel out. Using specific entropies $s_{\mathrm{c}}^{\mathrm{L}}(\tilde{T})$ and $s_{\mathrm{c}}^{\mathrm{G}}(\tilde{T})$, $S_{0}^{\mathrm{L}}=s_{\mathrm{c}}^{\mathrm{L}}(\tilde{T})N^{\rm L}_{0}$ and $S_{0}^{\mathrm{G}}=s_{\mathrm{c}}^{\mathrm{G}}(\tilde{T})N^{\rm G}_{0}$. The sum $F(\tilde{T},V^{\rm L}_{0},N^{\rm L}_{0})+F(\tilde{T},V^{\rm G}_{0},N^{\rm G}_{0})$ equals $F(\tilde{T},V,N)$. Applying these and substituting (B10) and (B11), (B12) is further transformed as
$F(\tilde{T},V,N)+\frac{\Xi}{2N}N^{\rm L}_{0}N^{\rm G}_{0}(s_{\mathrm{c}}^{\mathrm{L}}-s_{\mathrm{c}}^{\mathrm{G}}).$ (B13)
Applying the Clausius-Clapeyron relation, $s_{\mathrm{c}}^{\mathrm{L}}-s_{\mathrm{c}}^{\mathrm{G}}=(dp_{s}/dT)_{\tilde{T}}(v_{\mathrm{c}}^{\mathrm{L}}-v_{\mathrm{c}}^{\mathrm{G}})$ with $v_{\mathrm{c}}^{\mathrm{L}}=V^{\rm L}_{0}/N^{\rm L}_{0}$ and $v_{\mathrm{c}}^{\mathrm{G}}=V^{\rm G}_{0}/N^{\rm G}_{0}$, the second term becomes
$\frac{\Xi}{2N}\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}(V^{\rm L}_{0}N^{\rm G}_{0}-V^{\rm G}_{0}N^{\rm L}_{0}).$ (B14)
Using $V^{\rm L}_{0}N^{\rm G}_{0}-V^{\rm G}_{0}N^{\rm L}_{0}=-(VN^{\rm L}_{0}-NV^{\rm L}_{0})$ and $\Psi(N^{\rm L}_{0},V^{\rm L}_{0})=\frac{1}{2V}(N^{\rm L}_{0}V-NV^{\rm L}_{0})$, (B14) is further transformed into
$-\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{V}{N}\Psi(N^{\rm L}_{0},V^{\rm L}_{0}).$ (B15)
@@
${\cal F}_{g}(N^{\rm L}_{\rm s},V^{\rm L}_{\rm s})=F(\tilde{T},V,N)$
$\qquad-\Psi(N^{\rm L}_{0},V^{\rm L}_{0})mgL-\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{V}{N}\Psi(N^{\rm L}_{0},V^{\rm L}_{0})+O(\varepsilon^{2})$
$=F(\tilde{T},V,N)-\Psi(N^{\rm L}_{0},V^{\rm L}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2}),$
which is Eq. (19) in the main text (upon identifying $V/N=1/\bar{\rho}$ and neglecting $O(\varepsilon^{2})$).
A similar procedure for the other stationary solution $(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})$ (gas phase in the lower region), or by noting the symmetry $\Psi(N^{\rm G}_{0},V^{\rm G}_{0})=-\Psi(N^{\rm L}_{0},V^{\rm L}_{0})$ (where $N^{\rm G}_{0}=N-N^{\rm L}_{0}$, $V^{\rm G}_{0}=V-V^{\rm L}_{0}$), yields Eq. (20):
${\cal F}_{g}(N^{\rm G}_{\rm s},V^{\rm G}_{\rm s})$
$=F(\tilde{T},V,N)-\Psi(N^{\rm G}_{0},V^{\rm G}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2})$
$=F(\tilde{T},V,N)+\Psi(N^{\rm L}_{0},V^{\rm L}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg{|}_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2}).$

### v2

@@ Abstract
red In the linear response regime, we assume that local quantities exhibit linear profiles within each phase. Therefore, each state can be represented by its mean values. This allows us to map them to effective equilibrium systems, $(\tilde{T}^{\ell},\mathcal{V}^{\ell},\mathcal{N}^{\ell})$ and $(\tilde{T}^{\mathrm{u}},\mathcal{V}^{\mathrm{u}},\mathcal{N}^{\mathrm{u}})$, for which the following thermodynamic relations hold,
$dF^{\ell}=-S^{\ell}d\tilde{T}^{\ell}-\bar{p}^{\ell}d\mathcal{V}^{\ell}+\mu^{\ell}d\mathcal{N}^{\ell},$ (A1)
$dF^{\mathrm{u}}=-S^{\mathrm{u}}d\tilde{T}^{\mathrm{u}}-\bar{p}^{\mathrm{u}}d\mathcal{V}^{\mathrm{u}}+\mu^{\mathrm{u}}d\mathcal{N}^{\mathrm{u}},$ (A2)
[…]
Because $\tilde{T}^{\ell}$ and $\tilde{T}^{\mathrm{u}}$ are independent of $\mathcal{V}^{\ell}$ (see (13)), substituting the derivatives of the variational function (12) into the first equation of (14) yields
$\bar{p}^{\ell}-\bar{p}^{\mathrm{u}}=\frac{NmgL}{2V}$ (A3)
using the relation $\partial\Psi/\partial{\mathcal{V}^{\ell}}=-N/(2V)$ from (9). This relation (A3) reflects the overall force balance $(p_{1}-p_{2})A=Nmg$ and implies pressure continuity at the interface.
Similarly, the second equation of (14) yields
$\mu^{\ell}-\mu^{\mathrm{u}}-(S^{\ell}+S^{\mathrm{u}})\frac{\Xi}{2N}=\frac{mgL}{2}.$ (A4)
Using $T_{\mathrm{s}}$ defined in (16) of the main text, the global temperatures for each phase, given by (13), can be written as
$\tilde{T}^{\ell}=T_{\mathrm{s}}-\frac{\Xi}{2}\frac{\mathcal{N}^{\ell}}{N},\quad\tilde{T}^{\mathrm{u}}=T_{\mathrm{s}}+\frac{\Xi}{2}\frac{N-\mathcal{N}^{\ell}}{N}.$ (A5)
Furthermore, the mean pressures $\bar{p}^{\ell}$ and $\bar{p}^{\mathrm{u}}$ relate to the interface pressure $p_{\theta}$ via the force balance relations for each phase:
$(\bar{p}^{\ell}-p_{\theta})A=\frac{\mathcal{N}^{\ell}mg}{2},\quad(\bar{p}^{\mathrm{u}}-p_{\theta})A=-\frac{(N-\mathcal{N}^{\ell})mg}{2}.$ (A6)
Expanding $\mu^{\ell}(\tilde{T}^{\ell},\bar{p}^{\ell})$ and $\mu^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},\bar{p}^{\mathrm{u}})$ around $(T_{\mathrm{s}},p_{\theta})$ up to $O(\varepsilon)$, using $\partial\mu/\partial T=-S/N$, $\partial\mu/\partial p=V/N$, $\mathcal{V}^{\ell}=Ax_{\theta}$, and $V-\mathcal{V}^{\ell}=A(L-x_{\theta})$ along with (A5) and (A6), we have
$\mu^{\ell}(\tilde{T}^{\ell},\bar{p}^{\ell})=\mu^{\ell}(T_{\mathrm{s}},p_{\theta})+\frac{\Xi}{2}\frac{S^{\ell}}{N}+\frac{mgx_{\theta}}{2}+O(\varepsilon^{2}),$ (A7)
$\mu^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},\bar{p}^{\mathrm{u}})=\mu^{\mathrm{u}}(T_{\mathrm{s}},p_{\theta})-\frac{\Xi}{2}\frac{S^{\mathrm{u}}}{N}-\frac{mg(L-x_{\theta})}{2}+O(\varepsilon^{2}).$ (A8)
Substituting (A7) and (A8) into (A4), and neglecting terms of $O(\varepsilon^{2})$, we obtain (15) of the main text: $\mu^{\ell}(T_{\mathrm{s}},p_{\theta})=\mu^{\mathrm{u}}(T_{\mathrm{s}},p_{\theta})$.
@@
Consider the solution $(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})$ where the liquid phase settles in the lower region. We write
$N^{\mathrm{L}}_{\mathrm{s}}=N^{\mathrm{L}}_{0}+\Delta N,\quad V^{\mathrm{L}}_{\mathrm{s}}=V^{\mathrm{L}}_{0}+\Delta V,$ (B9)
where $N^{\mathrm{L}}_{0}$ and $V^{\mathrm{L}}_{0}$ are the particle number and volume of the liquid phase at $g=0$ and $\Xi=0$ (determined by the lever rule), and $\Delta N,\Delta V$ are deviations of $O(\varepsilon)$. According to (12) in the main text, the variational function is
$\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})=F^{\ell}(\tilde{T}^{\ell},V^{\mathrm{L}}_{\mathrm{s}},N^{\mathrm{L}}_{\mathrm{s}})$
$~+F^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},V-V^{\mathrm{L}}_{\mathrm{s}},N-N^{\mathrm{L}}_{\mathrm{s}})-\Psi(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})mgL.$ (B10)
@@
$\Psi(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})=\frac{N}{2}\left(\frac{N^{\mathrm{L}}_{0}}{N}-\frac{V^{\mathrm{L}}_{0}}{V}\right)+O(\varepsilon),$ (B11)
$\tilde{T}^{\ell}=\tilde{T}-\frac{\Xi}{2}\frac{N^{\mathrm{G}}_{0}}{N}+O(\varepsilon^{2}),$ (B12)
$\tilde{T}^{\mathrm{u}}=\tilde{T}+\frac{\Xi}{2}\frac{N^{\mathrm{L}}_{0}}{N}+O(\varepsilon^{2}).$ (B13)
@@
$F^{\ell}(\tilde{T}^{\ell},V^{\mathrm{L}}_{\mathrm{s}},N^{\mathrm{L}}_{\mathrm{s}})+F^{\mathrm{u}}(\tilde{T}^{\mathrm{u}},V-V^{\mathrm{L}}_{\mathrm{s}},N-N^{\mathrm{L}}_{\mathrm{s}})$
$=F(\tilde{T},V^{\mathrm{L}}_{0},N^{\mathrm{L}}_{0})+F(\tilde{T},V^{\mathrm{G}}_{0},N^{\mathrm{G}}_{0})$
$\quad-S_{0}^{\mathrm{L}}(\tilde{T}^{\ell}-\tilde{T})-p_{0}^{\mathrm{L}}\Delta V+\mu_{0}^{\mathrm{L}}\Delta N$
@@
[…]
$F(\tilde{T},V,N)+\frac{\Xi}{2N}N^{\mathrm{L}}_{0}N^{\mathrm{G}}_{0}(s_{\mathrm{c}}^{\mathrm{L}}-s_{\mathrm{c}}^{\mathrm{G}}).$ (B15)
Applying the Clausius-Clapeyron relation, $s_{\mathrm{c}}^{\mathrm{L}}-s_{\mathrm{c}}^{\mathrm{G}}=(dp_{s}/dT)_{\tilde{T}}(v_{\mathrm{c}}^{\mathrm{L}}-v_{\mathrm{c}}^{\mathrm{G}})$ with $v_{\mathrm{c}}^{\mathrm{L}}=V^{\mathrm{L}}_{0}/N^{\mathrm{L}}_{0}$ and $v_{\mathrm{c}}^{\mathrm{G}}=V^{\mathrm{G}}_{0}/N^{\mathrm{G}}_{0}$, the second term becomes
$\frac{\Xi}{2N}\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}(V^{\mathrm{L}}_{0}N^{\mathrm{G}}_{0}-V^{\mathrm{G}}_{0}N^{\mathrm{L}}_{0}).$ (B16)
Using $V^{\mathrm{L}}_{0}N^{\mathrm{G}}_{0}-V^{\mathrm{G}}_{0}N^{\mathrm{L}}_{0}=-(VN^{\mathrm{L}}_{0}-NV^{\mathrm{L}}_{0})$ and $\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})=\frac{1}{2V}(N^{\mathrm{L}}_{0}V-NV^{\mathrm{L}}_{0})$, (B16) is further transformed into
$-\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{V}{N}\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0}).$ (B17)
@@
$\mathcal{F}_{g}(N^{\mathrm{L}}_{\mathrm{s}},V^{\mathrm{L}}_{\mathrm{s}})=F(\tilde{T},V,N)$
$\qquad-\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})mgL-\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{V}{N}\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})+O(\varepsilon^{2})$
$=F(\tilde{T},V,N)-\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2}),$
which is (19) in the main text (upon identifying $V/N=1/\bar{\rho}$ and neglecting $O(\varepsilon^{2})$).
A similar procedure for the other stationary solution $(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})$ (gas phase in the lower region), or by noting the symmetry $\Psi(N^{\mathrm{G}}_{0},V^{\mathrm{G}}_{0})=-\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})$ (where $N^{\mathrm{G}}_{0}=N-N^{\mathrm{L}}_{0}$, $V^{\mathrm{G}}_{0}=V-V^{\mathrm{L}}_{0}$), yields (20):
$\mathcal{F}_{g}(N^{\mathrm{G}}_{\mathrm{s}},V^{\mathrm{G}}_{\mathrm{s}})$
$=F(\tilde{T},V,N)-\Psi(N^{\mathrm{G}}_{0},V^{\mathrm{G}}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2})$
$=F(\tilde{T},V,N)+\Psi(N^{\mathrm{L}}_{0},V^{\mathrm{L}}_{0})\left(mgL+\Xi\frac{dp_{\mathrm{s}}}{dT}\bigg|_{\tilde{T}}\frac{V}{N}\right)+O(\varepsilon^{2}).$